# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v5

@@ 4.3 Lipschitz Continuity of ICLRNN
A function $f:X\rightarrow\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L$ (i.e., L-Lipschitz) if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$:
$\|f(x_{1})-f(x_{2})\|_{2}\leq L\|x_{1}-x_{2}\|_{2}$ (11)
@@
$L(f)=L(\bm{\mathbf{W}})=\|\bm{\mathbf{W}}\|=\sigma_{max}(\bm{\mathbf{W}})$ (12)
@@
where $g$ represents a general transformation applied to the weighted inputs, $\bm{\mathbf{W}}^{(x)}$ and $\bm{\mathbf{U}}^{(h)}$ are weight matrices, and $\bm{\mathbf{x}}_{t}$ and $\bm{\mathbf{h}}_{t-1}$ are the inputs. The terms $\bm{\mathbf{W}}^{(x)}\bm{\mathbf{x}}_{t}$ and $\bm{\mathbf{U}}^{(h)}\bm{\mathbf{h}}_{t-1}$ contribute independently to the change in $\bm{\mathbf{h}}_{t}$. The Lipschitz constant $L(\bm{\mathbf{h}})$ of the mapping $\bm{\mathbf{x}}_{t},\bm{\mathbf{h}}_{t-1}\rightarrow\bm{\mathbf{h}}_{t}$ satisfies:
$L(\bm{\mathbf{h}}_{t})\leq\max(L(\bm{\mathbf{W}}^{(x)}),L(\bm{\mathbf{U}}^{(h)}))$ (14)
where $L(\bm{\mathbf{W}}^{(x)})$ is the Lipschitz constant (or spectral norm) of the weight matrix $\bm{\mathbf{W}}^{(x)}$ and $L(\bm{\mathbf{U}}^{(h)})$ is the Lipschitz constant (or spectral norm) of the weight matrix $\bm{\mathbf{U}}^{(h)}$.
@@
Consider a set of functions, e.g., $f:X\rightarrow\mathbb{R}$ and $g:X\rightarrow\mathbb{R}$ with Lipschitz constants $L(f)$ and $L(g)$ respectively, by taking their composition, denoted as $h=f\circ g$, the Lipschitz constant $L(h)$ of the resultant function $h$ satisfies:
$L(h)\leq L(f)\times L(g)$ (15)
@@
Based on the previously stated lemmas, we present the following proposition regarding the upper bound for the Lipschitz constant of the ICLRNN.
@@
Consider the $L$-layer ICLRNN. The Lipschitz constant of the output $\bm{\mathbf{y}}_{t}$ is upper bounded by 1 if and only if all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ and all activation functions $g_{i}$ have a Lipschitz constant upper bounded by 1.
@@
By unrolling the recurrent operations over time $t$ in Eq. (7), the output of ICLRNN can be described as a function, omitting the bias term:
@@
Thus, the Lipschitz constant of an ICLRNN is upper bounded by the product of the individual Lipschitz constants (i.e., Lemma 1 and Lemma 2):
$L(\bm{\mathbf{y}}_{t})\leq L(g_{t}^{(y)})\times L(\bm{\mathbf{W}}^{(y)})\times L(g_{t}^{(h)})\times\max(L(\bm{\mathbf{W}}^{(x)}),L(\bm{\mathbf{U}}^{(h)})\times\cdots\times L(g_{1}^{(h)})\times\max(L(\bm{\mathbf{W}}^{(x)}),L(\bm{\mathbf{U}}^{(h)})))$ (17)
By ensuring that the Lipschitz constants of all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ in the ICLRNN are upper bounded by 1 through spectral normalization, and that all activation functions $g_{i}$, such as ReLU, have Lipschitz constants also bounded by 1, the Lipschitz constant of the final output of the ICLRNN is guaranteed to be upper bounded by 1. ∎
@@
Since the function approximated by RNNs is not analytically tractable, we estimate the Lipschitz constant $L$ of the neural network numerically, as follows:
$L\approx\max_{x,y\in\mathbb{D}}\frac{||f(x)-f(y)||}{||x-y||}$ (18)
@@ 4.4 Input Convexity of ICLRNN
A function $f:X\rightarrow\mathbb{R}$ is convex if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$, with $x_{1}\neq x_{2}$, and $\forall~{}\lambda\in(0,1)$:
@@
and it is strictly convex if the relation $\leq$ becomes $<$.
Next, we develop the following proposition to show the convexity of $L$-layer ICLRNN.
@@
Consider the $L$-layer ICLRNN. The output $\bm{\mathbf{y}}_{t}$ is input convex if and only if all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ are non-negative and all activation functions $g_{i}$ are convex and non-decreasing.
@@
With Propositions 1 and 2, we present the following theorem to demonstrate that the ICLRNN is both input convex and Lipschitz-constrained.
@@
Consider the $L$-layer ICLRNN. Each element of the output $\bm{\mathbf{y}}_{t}$ is a convex, non-decreasing function of the input $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ (or just $\bm{\mathbf{x}}_{t}$), with the Lipschitz constant upper bounded by 1, at each time step $t=1,2,\dots,t$, for all $\hat{\bm{\mathbf{x}}}_{t}\in D\times D$ in a convex feasible space, provided that the following conditions are met: (1) All weights are non-negative, and their largest singular value is at most 1 at each time step $t$; (2) All activation functions are convex, non-decreasing, and Lipschitz continuous (e.g., ReLU) at each time step $t$.
@@
Using Proposition 1 and Proposition 2, we can demonstrate that the Lipschitz constant of the ICLRNN is upper-bounded by 1 and that it is input-convex. The Lipschitz constant is bounded by 1 because the weight matrices are spectral-normalized. Furthermore, since the weights are non-negative and the activation functions are convex and non-decreasing, the overall function remains convex. The proof follows directly from the fact that spectral normalization preserves convexity, as it is a linear transformation. Therefore, the ICLRNN is both input-convex and Lipschitz-constrained. ∎
5 Modeling and Control on a Chemical Process
In this section, we showcase the performance of the ICLRNN in modeling an ODE system through an example of a continuous stirred tank reactor (CSTR).
5.1 CSTR Modeling
@@
This reaction facilitates the conversion of reactant $A$ into product $B$. The CSTR incorporates a heating jacket responsible for supplying or extracting heat at a rate $Q$. The dynamic model of the CSTR is characterized by the ensuing material and energy balance equations:
$\frac{dC_{A}}{dt}$ $=\frac{F}{V}(C_{A0}-C_{A})-k_{0}e^{\frac{-E}{RT}}C_{A}^{2}$ (20a)
$\frac{dT}{dt}$ $=\frac{F}{V_{L}}(T_{0}-T)+\frac{-\triangle H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}^{2}+\frac{Q}{\rho_{L}C_{p}V}$ (20b)
@@
where $C_{A}$ denotes the concentration of reactant $A$, $F$ denotes the volumetric flow rate, $V_{L}$ denotes the volume of the reactant, $C_{A0}$ denotes the inlet concentration of reactant $A$, $k_{0}$ denotes the pre-exponential constant, $E$ denotes the activation energy, $R$ denotes the ideal gas constant, $T$ denotes the temperature, $T_{0}$ denotes the inlet temperature, $\Delta H$ denotes the enthalpy of reaction, $\rho_{L}$ denotes the constant density of the reactant, $C_{p}$ denotes the heat capacity, and $Q$ denotes the heat input rate. Moreover, $\bm{\mathbf{x}}^{T}=[C_{A}-C_{As},T-T_{s}]$ denotes the system states, where $T-T_{s}$ and $C_{A}-C_{As}$ represent the deviations in the temperature of the reactor and the concentration of $A$ from their respective steady-state values. $\bm{\mathbf{u}}^{T}=[\Delta C_{A0},\Delta Q]$ denotes the control actions, where $\Delta Q=Q-Q_{s}$ and $\Delta C_{A0}=C_{A0}-C_{A0_{s}}$ denote the manipulated inputs within this system, representing the alterations in the heat input rate and the inlet concentration of reactant $A$, respectively. In summary, the inputs consist of $[T_{t}-T_{s},C_{A,t}-C_{As},\Delta Q_{t},\Delta C_{A0,t}]$ at the current time step $t$. The outputs entail the state trajectory $[T_{t+1:n}-T_{s},C_{A,t+1:n}-C_{As}]$ over the subsequent $n$ time steps. In this case, $n$ is set to $10$. We conduct open-loop simulations for the CSTR of Eq. (20) using explicit Euler method. These simulations aim to collect data that mimics the real-world data for neural network training purposes. The detailed parameter values are listed in Table 1.
Parameter Value Range
$T_{s}$ $402~{}K$
$C_{A_{s}}$ $1.95~{}kmol/m^{3}$
$F$ $5~{}m^{3}/hr$
$V_{L}$ $1~{}m^{3}$
$Q_{s}$ $0.0~{}kJ/hr$
$T_{0}$ $300~{}K$
$R$ $8.314~{}kJ/kmol~{}K$
$C_{A0_{s}}$ $4~{}kmol/m^{3}$
$\rho_{L}$ $1000~{}kg/m^{3}$
$C_{p}$ $0.231~{}kJ/kg~{}K$
$E$ $5\times 10^{4}~{}kJ/kmol$
$k_{0}$ $8.46\times 10^{6}~{}m^{3}/kmol~{}hr$
$\Delta H$ $-1.15\times 10^{4}~{}kJ/kmol$
Table 1: Parameter values for CSTRs
@@ 5.2 Modeling Performance
Table 2: Hyperparameters of neural network models for CSTR
@@
ICLRNN (Ours) ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 28,436
@@
From Figs. 1 and 2, it is evident that conventional recurrent units, such as standard RNN and LSTM, lack robustness against noise. The higher Lipschitz constant observed for LSTM compared to RNN under increasing noise levels may result from overfitting to noisy data, likely due to its advanced gating mechanisms. In contrast, LRNN, ICRNN, and ICLRNN demonstrate robustness across different noise levels, maintaining small Lipschitz constants (less than 1), meaning they exhibit 1-Lipschitz behavior across various noise levels. Although ICRNN does not theoretically satisfy the Lipschitz property, the convexity constraint effectively enhances its robustness to noise. This robustness likely arises because the convexity constraint simplifies the model’s learning capability, enabling it to learn non-convex functions (such as noisy data) through convex approximations. This simplification inherently helps prevent “overfitting” to noise during training.
@@
The advantage of ICLRNN over LRNN lies in its lower testing MSE, indicating higher accuracy in modeling the system (see Fig. 1), and its significantly lower FLOPs, reflecting reduced model complexity. Specifically, for a 2-layer architecture with 64 neurons per layer, LRNN requires 159,508 FLOPs, whereas ICLRNN only requires 28,436 FLOPs - approximately 5.6 times fewer (see Table 2). This demonstrates that a less complex model can achieve more accurate results.
@@
Similarly, ICLRNN offers a significant advantage over ICRNN in terms of FLOPs, representing a less complex model under the same neural architecture. For a 2-layer architecture with 64 neurons per layer, ICRNN requires 79,892 FLOPs, while ICLRNN only requires 28,436 FLOPs - approximately 2.8 times fewer (see Table 2). Additionally, ICRNN has a theoretical structural limitation in that its hypothesis space cannot be very large. Due to the introduction of additional trainable parameters in the standard RNN architecture, ICRNN performs well when the hypothesis space is small (e.g., 2 layers with less than 128 neurons per layer). However, as the hypothesis space increases (e.g., 2 layers with more than 128 neurons per layer), training becomes unstable (see Table 3). This instability results in the MSE becoming Not a Number (NaN)  [^3] In computing, NaN, which stands for “Not a Number”, is a numeric data type that represents an undefined or unrepresentable value. It can be used to represent: infinity, the result of dividing by zero, a missing value, and the square root of a negative number., rendering the model untrainable unless stabilization techniques like layer normalization or batch normalization are employed. However, such techniques should be avoided in ICNN training because they introduce nonlinear scaling, violating the input convexity property of the final model. In summary, as the hypothesis space grows (e.g., by increasing the number of neurons), ICRNN’s testing MSE rises and eventually becomes NaN, whereas ICLRNN maintains a small and stable testing MSE.
@@
Table 3: Testing MSE of ICRNN and ICLRNN with respect to increasing hypothesis space
@@
Model Size ICRNN ICLRNN ICRNN ICLRNN

### v6

@@ 4.3 Lipschitz Continuity of ICL-RNN
A function $f:\mathbb{X}\rightarrow\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L_{f}$ (i.e., $L_{f}$-Lipschitz) if and only if the following inequality holds $\forall~x_{1},x_{2}\in\mathbb{X}$:
$\|f(x_{1})-f(x_{2})\|_{2}\leq L_{f}\|x_{1}-x_{2}\|_{2}$ (11)
@@
$L_{f}=\|\bm{\mathbf{W}}\|=\sigma_{max}(\bm{\mathbf{W}})$ (12)
@@
where $g$ represents a general transformation applied to the weighted inputs, $\bm{\mathbf{W}}^{(x)}$ and $\bm{\mathbf{U}}^{(h)}$ are weight matrices, and $\bm{\mathbf{x}}_{t}$ and $\bm{\mathbf{h}}_{t-1}$ are the inputs. The terms $\bm{\mathbf{W}}^{(x)}\bm{\mathbf{x}}_{t}$ and $\bm{\mathbf{U}}^{(h)}\bm{\mathbf{h}}_{t-1}$ contribute independently to the change in $\bm{\mathbf{h}}_{t}$. The Lipschitz constant $L_{h_{t}}$ of the mapping $\bm{\mathbf{x}}_{t},\bm{\mathbf{h}}_{t-1}\rightarrow\bm{\mathbf{h}}_{t}$ satisfies:
$L_{h_{t}}\leq\max(\|\bm{\mathbf{W}}^{(x)}\|,\|\bm{\mathbf{U}}^{(h)}\|)\times L_{g}$ (14)
where $\|\bm{\mathbf{W}}^{(x)}\|$ equals the Lipschitz constant (or spectral norm) of the weight matrix $\bm{\mathbf{W}}^{(x)}$ and $\|\bm{\mathbf{U}}^{(h)}\|$ equals the Lipschitz constant (or spectral norm) of the weight matrix $\bm{\mathbf{U}}^{(h)}$.
@@
Consider a set of functions, e.g., $f:\mathbb{X}\rightarrow\mathbb{R}$ and $g:\mathbb{X}\rightarrow\mathbb{R}$ with Lipschitz constants $L_{f}$ and $L_{g}$ respectively, by taking their composition, denoted as $h=f\circ g$, the Lipschitz constant $L_{h}$ of the resultant function $h$ satisfies:
$L_{h}\leq L_{f}\times L_{g}$ (15)
@@
Based on the previously stated lemmas, we present the following proposition regarding the upper bound for the Lipschitz constant of the ICL-RNN.
@@
Consider the $L$-layer ICL-RNN. The Lipschitz constant of the output $\bm{\mathbf{y}}_{t}$ is upper bounded by 1 if and only if all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ and all activation functions $g_{i}$ have a Lipschitz constant upper bounded by 1.
@@
By unrolling the recurrent operations over time $t$ in Eq. (7), the output of ICL-RNN can be described as a function, omitting the bias term:
@@
Thus, the Lipschitz constant of an ICL-RNN is upper bounded by the product of the individual Lipschitz constants (i.e., Lemma 1 and Lemma 2):
$L_{y_{t}}\leq L_{g_{t}^{(y)}}\times\|\bm{\mathbf{W}}^{(y)}\|\times L_{g_{t}^{(h)}}\times\max(\|\bm{\mathbf{W}}^{(x)}\|,\|\bm{\mathbf{U}}^{(h)}\|\times\cdots\times L_{g_{1}^{(h)}}\times\max(\|\bm{\mathbf{W}}^{(x)}\|,\|\bm{\mathbf{U}}^{(h)}\|))$ (17)
By ensuring that the Lipschitz constants of all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ in the ICL-RNN are upper bounded by 1 through spectral normalization, and that all activation functions $g_{i}$, such as ReLU, have Lipschitz constants also bounded by 1, the Lipschitz constant of the final output of the ICL-RNN is guaranteed to be upper bounded by 1. ∎
@@
Since the function approximated by RNNs is not analytically tractable, we estimate the Lipschitz constant $L_{nn}$ of the neural network numerically, as follows:
[…]
@@ 4.4 Input Convexity of ICL-RNN
A function $f:\mathbb{X}\rightarrow\mathbb{R}$ is convex if and only if the following inequality holds $\forall~(x_{1},x_{2})\in\mathbb{X}$, with $x_{1}\neq x_{2}$, and $\forall~\lambda\in(0,1)$:
@@
and it is strictly convex if the relation $\leq$ becomes <<.
Next, we develop the following proposition to show the convexity of $L$-layer ICL-RNN.
@@
Consider the $L$-layer ICL-RNN. The output $\bm{\mathbf{y}}_{t}$ is input convex if and only if all weights $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{U}}^{(h)},\bm{\mathbf{W}}^{(y)}\}$ are non-negative and all activation functions $g_{i}$ are convex and non-decreasing.
@@
With Propositions 1 and 2, we present the following theorem to demonstrate that the ICL-RNN is both input convex and Lipschitz-constrained.
@@
Consider the $L$-layer ICL-RNN. Each element of the output $\bm{\mathbf{y}}_{t}$ is a convex, non-decreasing function of the input $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ (or just $\bm{\mathbf{x}}_{t}$), with the Lipschitz constant upper bounded by 1, at each time step $t=1,2,\dots,t$, for all $\hat{\bm{\mathbf{x}}}_{t}\in D\times D$ in a convex feasible space, provided that the following conditions are met: (1) All weights are non-negative, and their largest singular value is at most 1 at each time step $t$; (2) All activation functions are convex, non-decreasing, and Lipschitz continuous (e.g., ReLU) at each time step $t$.
@@
Using Proposition 1 and Proposition 2, we can demonstrate that the Lipschitz constant of the ICL-RNN is upper-bounded by 1 and that it is input-convex. The Lipschitz constant is bounded by 1 because the weight matrices are spectral-normalized. Furthermore, since the weights are non-negative and the activation functions are convex and non-decreasing, the overall function remains convex. The proof follows directly from the fact that spectral normalization preserves convexity, as it is a linear transformation. Therefore, the ICL-RNN is both input-convex and Lipschitz-constrained. ∎
5 Case Study 1: Application To a Chemical Process Example
5.1 Process Description
@@
In this section, we demonstrate the effectiveness of ICL-RNN in a chemical reactor example of a CSTR system, which incorporates a heating jacket responsible for supplying or extracting heat at a rate $Q$ and facilitates the conversion of reactant $A$ into product $B$. The dynamic model of the CSTR is characterized by the ensuing material and energy balance equations:
$\frac{dC_{A}}{dt}$ $=\frac{F}{V_{L}}(C_{A0}-C_{A})-k_{0}e^{\frac{-E}{RT}}C_{A}^{2}$ (20a)
$\frac{dT}{dt}$ $=\frac{F}{V_{L}}(T_{0}-T)+\frac{-\triangle H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}^{2}+\frac{Q}{\rho_{L}C_{p}V_{L}}$ (20b)
@@
where $C_{A}$ denotes the concentration of reactant $A$, $F$ denotes the volumetric flow rate, $V_{L}$ denotes the volume of the reacting liquid in the tank, $C_{A0}$ denotes the inlet concentration of reactant $A$, $k_{0}$ denotes the pre-exponential constant, $E$ denotes the activation energy, $R$ denotes the ideal gas constant, $T$ denotes the temperature, $T_{0}$ denotes the inlet temperature, $\Delta H$ denotes the enthalpy of reaction, $\rho_{L}$ denotes the constant density of the reactant, $C_{p}$ denotes the heat capacity, and $Q$ denotes the heat input rate. Moreover, $\bm{\mathbf{x}}^{T}=[C_{A}-C_{As},T-T_{s}]$ denotes the system states, where $T-T_{s}$ and $C_{A}-C_{As}$ represent the deviations in the temperature of the reactor and the concentration of $A$ from their respective steady-state values. $\bm{\mathbf{u}}^{T}=[\Delta C_{A0},\Delta Q]$ denotes the control actions, where $\Delta Q=Q-Q_{s}$ and $\Delta C_{A0}=C_{A0}-C_{A0_{s}}$ denote the manipulated inputs within this system, representing the alterations in the heat input rate and the inlet concentration of reactant $A$, respectively. In summary, the inputs consist of $T_{t}-T_{s},C_{A,t}-C_{As},\Delta Q_{t},\Delta C_{A0,t}]$ at the current time step $t$. The outputs entail the state trajectory $T_{t+1:n}-T_{s},C_{A,t+1:n}-C_{As}]$ over the subsequent $n$ time steps. In this case, $n$ is set to $10$. We conduct open-loop simulations for the CSTR of Eq. (20) using explicit Euler method. These simulations aim to collect data that mimics the real-world data for neural network training purposes. The detailed parameter values are listed in Table 1.
Table 1: Parameter values for the CSTR model
Parameter Value Parameter Value
$T_{s}$ $402~\mathrm{K}$ $C_{A_{s}}$ $1.95~\mathrm{kmol/m^{3}}$
$F$ $5~\mathrm{m^{3}/hr}$ $V_{L}$ $1~\mathrm{m^{3}}$
$Q_{s}$ $0.0~\mathrm{kJ/hr}$ $T_{0}$ $300~\mathrm{K}$
$R$ $8.314~\mathrm{kJ/(kmol\cdot K)}$ $C_{A0_{s}}$ $4~\mathrm{kmol/m^{3}}$
$\rho_{L}$ $1000~\mathrm{kg/m^{3}}$ $C_{p}$ $0.231~\mathrm{kJ/(kg\cdot K)}$
$E$ $5\times 10^{4}~\mathrm{kJ/kmol}$ $k_{0}$ $8.46\times 10^{6}~\mathrm{m^{3}/(kmol\cdot hr)}$
$\Delta H$ $1.15\times 10^{4}~\mathrm{kJ/kmol}$
@@ 5.2 Modeling Performance
Table 2: Hyperparameters of neural network models
@@
ICL-RNN (Ours) ReLU $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}$ 28,436
@@
From Figs. 1 and 2, it is shown that conventional recurrent units, such as standard RNN and LSTM, lack robustness against noise. The higher Lipschitz constant observed for LSTM compared to RNN under increasing noise levels may result from overfitting to noisy data due to its advanced gating mechanisms. In contrast, LRNN, ICRNN, and ICL-RNN demonstrate robustness across different noise levels, maintaining small Lipschitz constants (less than 1), which implies that they exhibit 1-Lipschitz behavior across various noise levels. Although ICRNN does not theoretically satisfy the Lipschitz property, the convexity constraint effectively enhances its robustness to noise. This robustness is because the convexity constraint simplifies the model’s learning capability, enabling it to learn non-convex functions (such as noisy data) through convex approximations. This simplification inherently helps prevent “overfitting” to noise during training.
@@
The advantage of ICL-RNN over LRNN lies in its lower testing MSE, indicating higher accuracy in modeling the system (see Fig. 1), and its significantly lower FLOPs, reflecting reduced model complexity. Specifically, for a 2-layer architecture with 64 neurons per layer, LRNN requires 159,508 FLOPs, whereas ICL-RNN only requires 28,436 FLOPs - approximately 5.6 times fewer (see Table 2). This demonstrates that a less complex model can achieve more accurate results.
@@
Similarly, ICL-RNN offers a significant advantage over ICRNN in terms of FLOPs, representing a less complex model under the same neural architecture. For a 2-layer architecture with 64 neurons per layer, ICRNN requires 79,892 FLOPs, while ICL-RNN only requires 28,436 FLOPs - approximately 2.8 times fewer (see Table 2). Additionally, ICRNN has a theoretical structural limitation in that its hypothesis space cannot be very large. Due to the introduction of additional trainable parameters in the standard RNN architecture, ICRNN performs well when the hypothesis space is small (e.g., 2 layers with fewer than 128 neurons per layer). However, as the hypothesis space increases (e.g., 2 layers with more than 128 neurons per layer), training becomes unstable (see Table 3). This instability results in the MSE becoming Not a Number (NaN)  [^3] In computing, NaN, which stands for “Not a Number”, is a numeric data type that represents an undefined or unrepresentable value. It can be used to represent: infinity, the result of dividing by zero, a missing value, and the square root of a negative number., rendering the model untrainable unless stabilization techniques such as layer normalization or batch normalization are employed. However, such techniques should be avoided in ICNN training because they introduce nonlinear scaling, violating the input convexity property of the final model. In summary, as the hypothesis space grows (e.g., by increasing the number of neurons), ICRNN’s testing MSE rises and eventually becomes NaN, whereas ICL-RNN maintains a small and stable testing MSE.
@@
Table 3: Testing MSE of ICRNN and ICL-RNN with respect to increasing hypothesis space
@@
Model Size ICRNN ICL-RNN ICRNN ICL-RNN